### v1

# E-scooter effects on public transport demand: a case study in Santiago, Chile

Abstract
As cities adopt sustainable mobility solutions, electric scooters (e-scooters) offer both challenges and opportunities for public transportation systems. This study, the first in Latin America, examines the effects of e-scooters on public transport demand in Santiago, Chile, focusing on two scenarios: “generation” of trips (trips starting in study zones) and “attraction” of trips (trips ending in study zones). A negative binomial regression model was applied to data from public transport smart cards and e-scooter GPS. The methodology included urban area clustering and a differences-in-differences approach. The findings reveal significant regional differences: in the Central Region, public transport trips decreased by 21.38% in the generation scenario, while bus trips increased by 76.39%. In the Intermediate Region, metro trips increased by 70.05%, and in the Peripheral Region, bus trips increased by 84.64%. These results suggest that e-scooters reduce public transport usage in highly accessible areas but increase it in less accessible regions.
keywords:
Electric scooter , e-scooters , Micro-mobility , Mobility as a Service , Public Transport , Public Transit
journal: Transportation Research Part D
[…]
[aff1]organization=Universidad del Desarrollo, addressline=Av. Plaza 680, Las Condes, city=Santiago, country=Chile
[…]
[aff2]organization=Department of Computer Science, Universidad de Chile, addressline=Beauchef 851, city=Santiago, country=Chile
[…]
[aff3]organization=National Center for Artificial Intelligence (CENIA), addressline=Vicuna Mackenna 4860, Macul, city=Santiago, country=Chile
[…]
[aff4]organization=Department of Engineering Technology and Didactics, DTU Engineering Technology, city=Ballerup, country=Denmark
1 Introduction
Micromobility refers to modes of transportation utilizing small, lightweight vehicles powered by human effort or electricity, operated at low speeds (Dediu, 2019). These vehicles include bike-sharing systems, both docked and dockless e-scooters, as well as other vehicles weighing less than 500 kg, designed for individual travel in urban environments. Micro-mobility works well in dense areas (Tuli et al., 2021) and can help reduce traffic, cut emissions, and improve local economies.
Among these micro-mobility solutions, shared e-scooters have emerged as a prominent option. These vehicles typically operate through a Mobility as a Service (MaaS) application (Narayanan and Antoniou, 2023), allowing people to use their mobile phones to locate nearby scooters, reserve one for their trip, and ride to their destination. Depending on the service, e-scooters can sometimes be left parked anywhere, while in other cases they must be parked in specific designated spots. These services are usually led by private companies rather than transportation authorities and have been marketed as complements or enablers of public transportation systems, primarily as a first- or last-mile transportation mode (Zuniga-Garcia et al., 2022).
Nevertheless, the full extent of e-scooters’ impact on urban environments remains insufficiently understood. This is due to their recent introduction, the frequent lack of adequate regulation, and the limited availability of data (Aarhaug et al., 2023). On one hand, they offer the potential for modal shifts, complementing public transportation systems  (Latinopoulos et al., 2021; Zuniga-Garcia et al., 2022) and providing first- and last-mile solutions. On the other hand, they present challenges such as the displacement of unused vehicles, unauthorized use of public spaces, safety concerns for both riders and pedestrians (Stigson et al., 2021), conflicts with pedestrians (Gössling, 2020), and environmental issues related to their disposal (Echeverría-Su et al., 2023).
The scarcity of comprehensive studies on the overall impact of e-scooters complicates our understanding of their effects on urban environments, especially in relation to other transportation modes. Most existing research on this topic focuses on cities in the United States and Europe, which may not fully capture the complex interactions between e-scooters and public transport systems in other regions. Notably, in South America, there is a significant research gap in this field (Mitropoulos et al., 2023). This gap is particularly critical in cities like Santiago, where recent studies have shown significant changes in transportation patterns, including a reduction in mass-transit usage and an increase in motorized transport and taxi usage (Graells-Garrido et al., 2023a).
In this context, our study aims to address this critical gap by examining the effect of MaaS e-scooter availability on public transport usage in Santiago, Chile. Using a differences-in-differences methodology, we analyze official public transport data alongside e-scooter trip information from 2019, the peak usage period for these vehicles in Santiago. This is the first study of its kind in Latin America, offering empirical evidence on the interaction between e-scooters and public transport in a regional context. Our analysis provides valuable insights for urban planners and policymakers seeking to develop more effective and sustainable urban mobility strategies. By focusing on both trip generation and attraction scenarios, we capture the varying effects of e-scooter usage across different urban regions. Additionally, we apply a novel clustering method to automatically determine treatment and control zones, ensuring robust comparisons across different urban contexts. Furthermore, this study expands the scope of prior research by examining a multimodal public transport network, rather than focusing on a single mode, providing a comprehensive understanding of the interactions between e-scooters and public transport.
The paper structure is as follows: the next section presents a summary of the existing literature on the impact of shared e-scooters on public transportation usage. Following this, the study context and data description are presented. The methodology section is then detailed, divided into the study variables, clustering of zones of study, and the application of the differences-in-differences approach for the generation and attraction travel scenarios. This is followed by the results and discussion for these scenarios. Next, the paper addresses limitations and explores areas for improvement and future research directions. Finally, the paper concludes with a summary of the findings and their implications for urban mobility strategies
2 Literature Review
Most studies using e-scooter trips rely on spatial analysis and perform spatial and statistical analyses of trip data to estimate the relationship between e-scooter and public transportation trips (Vinagre Díaz et al., 2023; Yan et al., 2021; Aarhaug et al., 2023; Bai and Jiao, 2020; Jiao and Bai, 2020; Hosseinzadeh et al., 2021). Conversely, few studies have employed a quasi-experimental setting to understand this relationship (Ziedan et al., 2021a, b; Luo et al., 2021; Tyndall, 2022; Weschke et al., 2022).
Several spatial analysis studies have classified e-scooter trips as either complementary or substitutive to the transit system based on the proximity of e-scooter trip origins and destinations to transit stations (Vinagre Díaz et al., 2023; Yan et al., 2021; Aarhaug et al., 2023; Bai and Jiao, 2020; Hosseinzadeh et al., 2021; Jiao and Bai, 2020). For instance, Vinagre Díaz et al. (2023), based on data from Rome, reported that 50% of e-scooter trips either replace public transport or serve areas with inadequate public transport infrastructure. Similarly, Yan et al. (2021), using data from Washington, DC, found that approximately 70% of e-scooter trips could be considered substitutes, while around 20% were complementary to the transit system. Aarhaug et al. (2023) highlighted that e-scooters can simultaneously act as competitors and complements to public transport, based on their research conducted in Norwegian cities. Bai and Jiao (2020) utilized regression-based models to show that e-scooters can serve both roles in public transport systems in Austin and Minneapolis. Additionally, Hosseinzadeh et al. (2021) demonstrated a positive correlation between e-scooter usage and the density of bus stops in Louisville, indicating that e-scooters can complement public transport infrastructure, while Jiao and Bai (2020) confirmed that proximity to public transport stations is linked to increased e-scooter usage in Austin.
Regarding quasi-experimental analysis, there are two groups of studies. One group focuses on the effect of scooter trips on the number of bus riders  (Luo et al., 2021; Ziedan et al., 2021a, b), while the second group evaluates the demand for e-scooter trips based on light rail systems  (Tyndall, 2022; Weschke, 2023). Luo et al. (2021) conducted a spatial-temporal analysis using difference-in-difference models to assess the impacts of shared e-scooters on bus usage in Indianapolis, finding that 27% of e-scooter trips potentially competed with buses, while 29% were likely to increase bus usage. Ziedan et al. (2021a) used fixed-effects regression to investigate the effects on bus usage in Louisville, noting that shared e-scooters did not significantly impact local bus usage. Conversely, Ziedan et al. (2021b) classified trips by trip purpose using data from Nashville, Tennessee. The study found that routine weekday trips are related to a 0.94% decrease in bus usage, while social e-scooter trips, are related to a 0.86% increase in weekday bus usage.
Tyndall (2022) analyzes e-scooter trip data in Seattle, while Weschke (2023) uses data from Berlin. Tyndall (2022) explores the expansion of a light rail system in the U.S. as a natural experiment, applying a difference-in-difference approach. The study finds that the introduction of a new light rail station led to a significant increase in the presence of shared bicycles and e-scooters within a five-minute walking distance from the station. Similarly, Weschke (2023) examines the extension of a metro line in Berlin, Germany, through a quasi-experiment, revealing that e-scooter demand increases at newly opened stations, while it significantly decreases when a station is closed. Both studies concluded that public transport acts as a complement for shared e-scooters.
Table 1 below provides a summary of the key studies that examine the relationship between e-scooter trips and public transportation, highlighting their methodologies, data, and main findings.
Authors Method Data Place Main result
Vinagre Díaz et al. (2023) Clustering technique E-scooter trips, and rail-based transit network Rome, Italy 50% of scooter trips either replace transit or cover areas with low transit accessibility
Yan et al. (2021) Spatiotemporal analysis E-scooter trips, bikeshare network, and transit network Washington, DC, US 20% of trips are last-mile connections to transit and 70% are transit substitutes
Ziedan et al. (2021a) Fixed effects regression models E-scooter trips, network and trips bus ridership data Louisville, US Complementary effect on express bus route - transit increase 5 trips for every 10 e-scooter trips
Luo et al. (2021) Spatial-temporal analysis with a Difference-in-Differences model E-scooter trips, network and trips bus ridership data Indianapolis, Indiana, US 27% of e-scooter trips compete with bus trips and 29% complement the bus system
Ziedan et al. (2021b) Fixed effects regression models E-scooter trips, and trips bus ridership data Nashville, Tennessee, US Net impact of e-scooter trips is -0.08% on bus ridership
Aarhaug et al. (2023) Logit regression model E-scooter trips, travel survey, planning transport data Oslo, Norway 50.5% of e-scooter trips compete with public transport and 20% complement public transport
Baek et al. (2021) Mixed logit model Stated Preference survey Seoul, South Korea 40.6% of trips indicating a preference for e-scooters, 39.2% for buses, and 20.2% for walking
Bai and Jiao (2020) Spatial analysis and negative binomial regression models E-scooter trips, build environment, and demographic data Austin, Texas, US; Minneapolis, Minnesota, US 52% and 47% (Austin and Minneapolis) of e-scooter trips compete with public transport and 31% and 29% of them complement public transport
Jiao and Bai (2020) Spatial analysis and a negative binomial regression model E-scooter trips, build environment, and demographic data Austin, Texas, US 52% of e-scooter trips compete with public transport and 31% of them complement public transport
Hosseinzadeh et al. (2021) Geographically Weighted Regression E-scooter trips, build environment, and demographic data Louisville, Kentucky, US 50% of e-scooter trips substituting public transportation and 25% complementing it
Tyndall (2022) Difference-in-difference E-scooter and e-bikes parking locations, transit network, demographic, and geographic data Seattle, Washington, US The flow of new micromobility vehicles increased significantly within a 5 min walking radius of the new rail station
Weschke (2023) Difference-in-differences E-scooter trips, OpenStreetMap Data Berlin, Germany Public transport acts as complement for shared e-scooters
Table 1: Summary of Studies on the Impact of E-Scooter Trips on Public Transportation. This table summarizes studies on the impact of e-scooter trips on public transportation. The columns include the authors and year of publication, the methodological approach (e.g., clustering, spatiotemporal analysis, regression models), the types of data used (e.g., e-scooter trips, network data, bus ridership, travel surveys), the geographic location of the study, and the main findings regarding how e-scooter trips affect public transportation.
These previous studies have laid the foundation for understanding the interaction between e-scooters and public transport. Building on their insights, our study expands prior research (Luo et al., 2021; Ziedan et al., 2021a, b; Tyndall, 2022; Weschke, 2023) by employing a quasi-experimental approach. The key differences and contributions of our study are:
  • 1.
    Our study is conducted in Santiago, Chile, offering a Latin American perspective that is underrepresented in the existing literature. This regional focus is crucial, given the unique urban dynamics and public transport challenges in Latin American cities.
  • 2.
    We consider a multimodal public transport network, whereas Ziedan et al. (2021a, b), Luo et al. (2021), Tyndall (2022), and Weschke (2023) focus solely on specific transit modes.
  • 3.
    Our study encompasses a large-scale transport network, covering the entire city of Santiago, including central, intermediate, and peripheral regions. This comprehensive approach allows us to capture a wider range of interactions between e-scooters and public transport across different urban settings.
  • 4.
    We propose a geographical and transport of mode analysis to gain a deeper understanding of which public transport trips are related to e-scooter trips.
  • 5.
    Our study proposes a clustering methodology that automatically determines control and treatment zones, as opposed to other studies that select these zones manually.
3 Data Description
This section describes the data sources used in our study of transportation demand in the Great Metropolitan Area of Santiago, Chile. We analyze both public transport and shared e-scooter usage across two distinct time periods, utilizing comprehensive datasets to examine the impact of e-scooters on public transport demand.
Santiago, Chile’s capital and largest city, had a population of nearly 8 million and covered an urban area of 867.75 square kilometers as of the 2017 Census. The city comprises over 40 independent municipalities, with 35 in the urban core and the rest on the periphery (see Fig. 1 a). Santiago’s public transportation system is integrated and multimodal, featuring a nearly flat fare structure across metro, urban buses, and a rail service. This system allows up to three transfers within a two-hour window. The bip! smart card is the sole payment method accepted, requiring validation when boarding buses or entering metro stations, but not upon exiting. Consequently, the smart card data only records trip origins. In addition to public transport, Santiago offers extensive taxi services, several ride-hailing applications, and a public bike-sharing system.
For this work, we analyze two types of transportation demand: public transport and shared e-scooters. Our analysis focuses on two distinct time periods:
  1. 1.
    Public transport data: We examined bip! records from the second week of May 2018 and the second week of May 2019. These weeks were chosen as they are similar in terms of typical urban activity, without any particular events or interventions in the city. It is important to note that by 2019, the newly inaugurated Metro Line 3 was already in operation, which we considered in our analysis.
  2. 2.
    E-scooter data: We analyzed trips from a shared e-scooter application, covering the period from October 10th, 2018 to April 15th, 2019. This dataset includes the period when e-scooter services were introduced and became established in Santiago.
We selected these specific time periods to capture both the typical urban activity patterns and the introduction and establishment of e-scooter services in Santiago. The inclusion of data from 2018 and 2019 also allows us to account for the impact of the newly inaugurated Metro Line 3, which began operations between these two periods.
In this study, the unit of analysis is the geographical zoning of the Metropolitan Public Transport Directory (DTPM). It consists of 804 transportation analysis zones (see Fig. 1 b). These zones have an average area of 1.23 Km2. Next, we describe the datasets used to measure the effect of e-scooters in public transport demand.
Figure 1: Area under analysis. Left (a): Administrative boundaries of municipalities (orange) and of Traffic Analysis Zones (purple). Right (b): Metro and train lines operating in the city. The red line (Metro Line 3) was inaugurated in late 2018, between our periods under study. OpenStreetMap data and Map tiles © CARTO are used under a Creative Commons License.
Our primary dataset originates from the ADATRAP system (Gschwender et al., 2016), provided by DTPM. ADATRAP consolidates multiple data sources from Santiago’s public transport system, including smart card transactions, bus GPS records, and vehicle operator information, among others. These data are processed and enhanced by analyzing user travel patterns and the spatial distribution of distances to boarding points (Munizaga and Palma, 2012). Trips are reported at bus stop or metro station level, which can be aggregated into zones (see Fig. 2.a to view the spatial distribution of all trips under study), municipalities or macro-areas that contain multiple municipalities.
Figure 2: Distributions of Public Transport trips. Left (a): a choropleth map of total trips with respect to Traffic Analysis Zones. Top right (b): total trips per day, per year. Bottom right (c): relative distribution of trip starting periods for each day, where the periods are defined in Table A.6.
Through an agreement between a e-scooter operator and the Las Condes municipality in Santiago (see Fig. 4), we were able to analyze the origins and destinations of e-scooter trips in Las Condes. The e-scooter operator provided a dataset of GPS traces. Trips in this dataset were detected using a stop-based algorithm (Graells-Garrido and Saez-Trumper, 2016). Each observation indicated whether the vehicle was being under control of a traveler or a staff person, as such, we were able to discard relocation trips. In total, this dataset contains 45K trips between October 10th, 2018, and April 15th, 2019. The time of the day and distance distributions of those trips (and ADATRAP trips for comparison) are shown in Fig. 3. Notably, e-scooter trips have a less pronounced morning peak (6:00-9:00 hrs) in comparison to public transport trips, although the afternoon peak (5:30-20:30 hrs) shows a similar trend. At lunch time (12:00-14:00 hrs), e-scooter trips present a more pronounced distribution than public transport, which is arguably related to the area of availability, as Las Condes is a municipality that concentrates high-density working areas and business centers. Notably, trips in e-scooters are shorter than public transportation trips (with means of 1.44 Km and 9.25 Km, respectively). The origins and destinations of trips seem to be closely related to the metro stations available in their area of operation (see Fig. 4).
Figure 3: Distributions of trip features from Public Transport and shared e-scooters. Top (a): Proportion of trips with respect to hour of the day, per year. Bottom left (b): Histogram of trip distances in e-scooter trips. Bottom right (c): Histogram of trip distances in Public Transport.
Figure 4: Origins and destinations of shared e-scooter trips. The top row shows trip origins in three periods of the day (morning peak, lunch and afternoon peak), whereas the bottom row shows trip destinations for the same period. The map on the right highlights the area of availability of the e-scooter service: the municipality of Las Condes.
Finally, we utilized demographic data from the 2017 Chilean Census  (National Institute of Statistics, 2017). The variables analyzed included population counts by age and sex, as well as the average educational level (measured in years of schooling). Educational level was used as a proxy for income to capture the socio-economic characteristics of each zone. As the census does not report individual income directly, we used years of schooling as a proxy for income. This approach is supported by research showing strong correlations between educational attainment, income distribution, and inequality (Gregorio and Lee, 2002; Jerrim and Macmillan, 2015). All variables were aggregated at the traffic analysis zone level. We have also used the official General Transit Feed Specification (GTFS) provided by DTPM to calculate some accessibility indicators. We describe these datasets in the next section, as their variables are intricately related to our methodology to quantify the effect of shared e-scooters on public transport demand.
4 Methodology
To examine the impact of scooters, we conducted a study utilizing the difference-in-differences (Diff-in-Diff) approach. Our spatial unit of analysis consisted of the geographical zones defined by the DTPM. The urban areas (geographical zones) were divided into three distinct groups using a clustering method to account for spatial heterogeneity and ensure that comparisons between control and treatment zones within each cluster were valid. In this section, we present the details of the variable selection process for clustering and the application of the Diff-in-Diff methodology.
Selection of Study Variables
We compiled an initial list of variables that are typically considered to influence travel behavior and assessed their relevance in the context of Santiago and this study. The selected variables include sociodemographic variables from the 2017 census, such as underage population, working population, retired population, percentage of female population, average educational level, and population density. Additionally, we included variables from ADATRAP dataset to capture travel behavior, such as the average number of public transport trips per person, average time of the first public transport trip made each day, and average public transport trip distance. We also included variables to determine each zone’s accessibility to public transport, mapping the metro network within Santiago and counting the number of metro stations and bus stops per zone using data from OpenStreetMap and the General Transit Feed Specification (GTFS). These indicators were calculated for each zone in our design. A comprehensive list of these variables can be found in Table 2. Additionally, Figure 5 and Figure 6 show the distribution of these variables across the different zones of the city. We can see that there’s a central point in the city that is highly densely populated, and as we get farther away from it the population spreads out.
Name description source use unit mean std
under age population Percentage of the population that’s younger than 18. 1 Zone clustering % 22.50 4.78
working population Percentage of the population that’s between the ages of 18 and 65. 1 Zone clustering % 65.60 4.44
retired population Percentage of the population that’s older than 65. 1 Zone clustering % 11.90 4.61
female population Percentage of the population that’s female. 1 Zone clustering % 51.50 2.11
avg. educational level Average of years of education received by people over 18. 1 Zone clustering years 11.86 1.80
bus stops Number of bus stops in the zone. 2 Zone clustering count 14.09 8.11
bus stop density Bus stops per square meter. 2 Zone clustering stops/km2 14.15 7.90
bus stops per 1000 people Bus stops per 1000 inhabitants. 1, 2 Zone clustering stops/person 2.50 4.13
population density Inhabitants per square meter. 1 Zone clustering person/km2 7696.17 4591.17
avg. number of trips per person Average number of public transport trips each person makes. 3 Zone clustering Regression count 1.91 0.28
avg. time of first trip Average hour of the first trip made on public transport each day. 3 Zone clustering hour 10.99 0.40
avg. trip distance Average distance of trips made on public transport originating in the zone. 3 Zone clustering meters 9218.33 3053.81
Table 2: Summary of Zone Clustering Variables. This table presents a comprehensive summary of the variables employed in our study. Each column includes the variable name, a brief description, the data source (represented by a number), the role of the variable in the analysis (use), its unit of measurement, and key statistical indicators such as the mean and standard deviation. The variables span demographic information (e.g., underage population, working population), educational level, and transport-related metrics (e.g., number of bus stops, bus stop density). The data sources are as follows: (1) 2017 Census, (2) OpenStreetMap, and (3) ADATRAP.
Figure 5: Spatial Distribution of Demographic Characteristics. This set of maps illustrates the spatial distribution of demographic indicators, including underage population, working population, retired population, female population, average educational level, and population density. These visualizations provide insights into the sociodemographic structure.
Figure 6: Spatial Distribution of Transportation and Mobility Metrics. This figure displays the spatial distribution of transportation infrastructure and usage metrics, such as the number of bus stops, bus stop density, bus stops per 100 people, average trips per person, average time of the first trip, and average trip distance.
Clustering and Definition of Treatment and Control Groups
In this study, we aimed to create meaningful spatial clusters that reflect both statistical similarities and geographic proximity. The process of creating regions, known as regionalization  (Duque et al., 2007), is crucial in this context. Regionalization is a specific form of clustering aimed at grouping observations that are alike in both their statistical characteristics and spatial proximity. This ensures that the resulting zones are not isolated but are similar to their neighboring areas, aligning with our goal of meaningful spatial clustering.
We started by using Moran’s I to analyze the spatial autocorrelation of variables from the previous section. This method helped identify variables with significant spatial patterns, ensuring that the selected variables were not randomly distributed across space but had meaningful spatial dependencies. A positive Moran’s I value indicates clustering over space, while a negative Moran’s I indicates dispersion. We chose 0.25 as a threshold for qualifying variables. Additionally, we established a selection criterion based on the p-sim value of each variable, representing the likelihood of the observed values being randomly spatially distributed. We set 0.05 as the upper limit for the p-sim value of each variable.
Based on these criteria, we selected the following cluster variables: percentage of underage population, percentage of working population, percentage of retired population, average educational level, percentage of female population, bus stop density, population density, average number of public transport trips made per person, average time of the first public transport trip of the day, and average public transport trip distance.
We implemented two clustering algorithms using Scikit-learn, a free, open-source Python library that provides tools for machine learning applications: k-means (Lloyd, 1982) and agglomerative hierarchical clustering. For agglomerative hierarchical clustering, we used the Ward method as the linkage criterion (Ward Jr, 1963), which minimizes the total within-cluster variance when merging clusters. We employed two different connectivity approaches for this method: a k-nearest neighbors (KNN) contiguity matrix and a sparse connectivity matrix. Each approach was applied across a range of 3 to 10 clusters.
Afterward, we used the Calinski-Harabasz score (Caliński and Harabasz, 1974) to compare the performance of these techniques and configurations to determine the best algorithm. The Calinski-Harabasz (CH) index, also known as the variance ratio criterion, measures the ratio of separation (the distance between cluster centroids) to cohesion (the distance between points within a cluster and their centroid). A higher score indicates better-defined clusters. As illustrated in Figure 7, the k-means algorithm with 3 clusters achieved the highest score, making it our chosen model.
Figure 7: This figure shows the Calinski-Harabasz (CH) scores versus the number of clusters for three clustering methods: k-means (purple), agglomerative hierarchical with KNN connectivity (blue), and agglomerative hierarchical with sparse connectivity (teal). The x-axis represents the number of clusters (2 to 10), and the y-axis shows the CH score.
The resulting regionalization (groups of zones) after applying the k-means clustering algorithm with k=3 is presented in Figure 8. The three clusters, named according to their location in the city, almost form concentric circles over Santiago. The zones in light yellow represent the central area (Central Region), the zones in purple represent the periphery (Peripheral Region), and the zones in turquoise correspond to an intermediate area of the city (Intermediate Region).
Figure 8: Spatial distribution of zones resulting from the k-means clustering algorithm with k=3. The zones in blue represent the Central Region, the zones in orange correspond to an intermediate area of the city (Intermediate Region), and the zones in purple represent the Peripheral Region.
After using the k-means algorithm to divide the city into three regions, we measured how the study variables were spread out in each region. This analysis revealed distinct characteristics for each region. Figure 9 and Figure 10 illustrate the characteristics of three distinct regions (Central, Peripheral, and Intermediate) in terms of demographic and transportation metrics. The Central Region, represented in blue, shows lower values in population categories (underage, working, and female populations) but higher educational levels and bus stop density. This region is densely populated and has a mostly adult population that is highly educated. Public transport trips originating in these areas are short, averaging less than a kilometer, and infrequent, with people making less than 2 trips per day on average. The Peripheral Region, in yellow, has a higher proportion of retired and female populations, alongside the highest population density and average trip distance. It also has the highest proportion of underage population and the lowest educational attainment among adults. Public transport trips starting in this zone are longer and more frequent than in the other two regions. The Intermediate Region, in purple, balances these metrics, displaying moderate values in all categories, suggesting a transitional nature between the Central and Peripheral regions. Each cluster’s transportation usage is distinct, with the central region showing the highest average number of trips per person and the shortest average trip distance. It is important to note that the density of bus stops, and therefore the level of accessibility to public transport, is higher in the central zone than in the intermediate and peripheral zones.
Figure 9: Demographic Characteristics Across Central, Peripheral, and Intermediate Regions. The figure shows box plots of demographic metrics for three regions: Central (blue), Peripheral (purple), and orange (purple).
Figure 10: Transportation Characteristics Across Central, Peripheral, and Intermediate Regions. The figure shows box plots of transportation metrics for three regions: Central (blue), Peripheral (orange), and Intermediate (purple).
After applying the clustering algorithm and dividing the zones into three groups of regions, we excluded 56 zones from the study despite them having no origin and destination trips involving scooters, as they could still experience mobility impacts post-intervention. The criteria for exclusion were zones intersecting with a 1,440 m buffer applied to each scooter trajectory, a distance chosen based on the average trip distance calculated from our data. Figure 11 displays the control zones, treatment zones, and the excluded zones. Figure 12, along with Table 3, presents the distribution and detailed breakdown of control and treatment zones by cluster, excluding the 56 mentioned zones.
Figure 11: Spatial Distribution of Scooter Trips, Treatment, and Control Zones. The dark bluelines represent scooter trips. Treatment zones, highlighted in light blue, are those where scooter trips originated or ended. Zones highlighted in yellow are treatment zones with a buffer of 1,440 meters, chosen based on the average trip distance calculated from our data. Control zones, shown in gray, are not directly impacted by scooter trips or their buffer zones.
Figure 12: Distribution of the treatment and control zones by region. The map displays the zones within the study area.The figure presents the spatial distribution of zones across central, intermediate, and peripheral regions, categorized as either treatment or control zones. Darker shades correspond to control areas, while lighter shades denote treatment areas, with white zones representing regions excluded from the analysis
Region Control Zones Treatment Zones
Central Region 53 5
Intermediate Region 304 57
Peripheral Region 312 9
Table 3: Number of Control and Treatment Zones by Region. The column Control Zones indicates the number of control zones, while Treatment Zones indicates the number of treatment zones, excluding the 56 zones that intersect with a 1,440 m buffer around scooter trajectories.
Regression Model
Finally, we measured how the implementation of e-scooters affects public transportation demand. Given the number of public transportation trips often exhibits significant variability and over-dispersion, meaning that the variance exceeds the mean. This makes traditional Poisson regression unsuitable for our dataset. Instead, we applied two distinct Negative Binomial NB regression models (Cameron and Trivedi, 1986): one for trips generation (trips originating from public transport stations) and one for trips attraction (trips ending at public transport stations). The NB models are well-suited to handle over-dispersed count data and allow us to better capture the variability in public transportation trips across different zones, providing a more accurate understanding of the factors influencing both trip generation and attraction.
The general form of the NB model is defined as follows:
$f(y_{i}\mid\mathbf{x}_{i})=\frac{\Gamma(y_{i}+\alpha)}{y_{i}!\Gamma(\alpha)}\left(\frac{\alpha}{\alpha+u_{i}}\right)^{\alpha}\left(\frac{u_{i}}{\alpha+u_{i}}\right)^{y_{i}},$
where $y_{i}$ is the response variable, i.e., the count of public transportation trips in Traffic Analysis Zones (TAZ), $\mathbf{x}_{i}$ is the set of independent variables, $\Gamma(\cdot)$ is the gamma function, and $\alpha$ is the dispersion parameter. The conditional mean and variance are given by:
$\mathbf{E}(y_{i}\mid\mathbf{x}_{i})=u_{i}=\exp(\mathbf{x}_{i}^{T}\mathbf{\beta});\text{Var}(y_{i}\mid\mathbf{x}_{i})=u_{i}\left(1+\frac{1}{\alpha}u_{i}\right).$
This choice of model is coherent with studies that have applied similar approaches to transportation demand across various modes  (e.g., Schimohr and Scheiner, 2021; Hu et al., 2022; Manville et al., 2023), and with previous research analyzing behavioral and spatial patterns in Santiago  (Graells-Garrido et al., 2017; Beiró et al., 2018; Graells-Garrido et al., 2023b).
Here, the set of independent variables that define each trip count contains the following:
  • 1.
    year: a binary variable which has two possible values: 2019 and 2018, with 2018 as reference value;
  • 2.
    type of day: a categorical variable with four possible values: monday-to-thursday, friday, saturday, and sunday. monday-to-thursday is used as the reference category.
  • 3.
    time period: a categorical variable that describes the time of the day when the trip was performed, including the following values: morning valley, morning peak, lunch, afternoon valley, afternoon peak, night, and night valley, with morning valley as reference value. The time periods are defined in the table provided in the Appendix A (see Table A.6);
  • 4.
    scooter trips origins: a dummy variable where 1 indicates that there are e-scooter trips originating in the specific TAZ, and 0 indicates none;
  • 5.
    scooter trips destinations: a dummy variable where 1 indicates that there are e-scooter trips ending in the specific TAZ, and 0 indicates none;
  • 6.
    working population: refers to the log-transformed count of the working population that resides in the specific TAZ;
  • 7.
    relative educational level: denotes the average educational level for the specific TAZ measured as years of education, relative to the regional average;
  • 8.
    metro: denotes the number of metro lines within a distance of less than 750 meters from the specific TAZ area.
Using these independent variables, we define the negative binomial regression models for both generation and attraction of trips. For the generation model, which examines trips originating from public transport stations, we use the following expression:
$\begin{split}\ln\left({\textbf{E}}[y_{\text{gen}}\mid\mathbf{x}_{\text{gen}}]\right)=&\beta_{0}+\beta_{1}\text{year}+\beta_{2}\text{type\_of\_day}+\beta_{3}\text{time\_period }+\\ &\beta_{4}\text{scooter\_trips\_destinations}+\beta_{5}\text{working\_population }+\\ &\beta_{6}\text{relative\_educational\_level}+\beta_{7}\text{metro}\end{split}$
Similarly, for the attraction model, which examines trips ending at public transport stations, the model is expressed as:
$\begin{split}\ln\left({\textbf{E}}[y_{\text{att}}\mid\mathbf{x}_{\text{att}}]\right)=&\beta_{0}+\beta_{1}\text{year}+\beta_{2}\text{type\_of\_day}+\beta_{3}\text{time\_period }+\\ &\beta_{4}\text{scooter\_trips\_origins}+\beta_{5}\text{working\_population }+\\ &\beta_{6}\text{relative\_educational\_level}+\beta_{7}\text{metro}\end{split}$ (1)
The key difference between these models is in the scooter trip variable: the generation model $f(y_{gen}\mid\mathbf{x}_{gen})$ uses scooter trip destinations, while the attraction model $f(y_{att}\mid\mathbf{x}_{att})$ uses scooter trip origins. This distinction allows us to capture the different effects of e-scooters on public transport usage in each scenario.
For each scenario (generation and attraction), we developed three separate models: 1) A model for bus trips only; 2) a model for metro trips only; and 3) a combined model for both bus and metro trips. In all models, the dependent variable is the count of public transport trips in each Traffic Analysis Zone (TAZ). We applied the regression to three regions (Central, Intermediate and Peripheral), each containing its respective control and treatment areas. Our analysis focused on two distinct time periods in fall: the second week of May 2018, and the second week of May 2019. A differences-in-differences approach was utilized to control for temporal trends and other unobservable differences between the zones that implemented scooters (treatment group) and those that did not (control group).
This approach captures the effects within each region individually, providing a more detailed understanding of the relationship between scooter implementation and public transportation usage across different zones and transport modes. The inclusion of specific days of the week and periods of the day in our models accounts for variations in transportation usage patterns due to different daily and temporal dynamics.
5 Results
In this section, we present the findings of our study on the impact of shared e-scooters (e-scooters) on public transportation usage in Santiago, Chile. Our analysis, using a negative binomial regression model and a difference-in-differences approach, examines how e-scooter trips affect public transport demand across different urban regions, considering both generation and attraction scenarios. For the metro trip generation and attraction models, we focused on the Central and Intermediate areas, as there are very few metro stations in the Peripheral Region.
Generation of Trips
Tables in Appendix C show the results of negative binomial models for public transport trip generation, and Table 4 shows the marginal effects of scooter destination zones, located within different regions, on public transport trip generation. In the Central Region, shared e-scooter trips are linked to fewer combined bus and metro trips, with an average reduction of 136.66 trips per hour per zone. This represents a 21.38% reduction, considering the region’s average of 639.26 combined bus and metro trips per hour. However, when focusing specifically on bus trips, shared e-scooters contribute to an increase of 145.22 additional bus trips per hour per zone, reflecting a 76.39% rise from the average of 190.09 bus trips per hour. For metro trips, the impact of e-scooters is not statistically significant.
In the Intermediate Region, which also comprises multiple zones, shared e-scooter trips contribute to an average of 99.77 additional bus and metro trips per hour per zone, leading to a 56.83% increase in combined public transport usage, considering the region’s average of 175.56 trips per hour. The impact on bus trips is more modest, with an increase of 20.10 additional bus trips per hour per zone, corresponding to a 22.52% rise from the region’s average of 89.26 bus trips per hour. For metro trips, e-scooters result in 365.30 additional trips per hour per zone, representing a 70.05% increase relative to the region’s average of 521.46 metro trips per hour.
In the Peripheral Region, shared e-scooter trips generate an average of 45.40 additional bus and metro trips per hour per zone, resulting in a 40.17% increase in combined public transport usage, considering the region’s average of 113.03 trips per hour. The effect on bus trips is more pronounced, with 72.82 additional bus trips per hour per zone, representing an 84.64% increase from the region’s average of 86.04 bus trips per hour.
Mode of Transport Category Central Intermediate Peripheral
Bus or Metro E-scooter impact on trips/hour -136.66 99.77 45.40
Average trips 639.26 175.56 113.03
% Change in trips -21.38% 56.83% 40.17%
Bus E-scooter impact on trips/hour 145.22 20.10 72.82
Average trips 190.09 89.26 86.04
% Change in trips 76.39% 22.52% 84.64%
Metro E-scooter impact on trips/hour - 365.30 -
Average trips - 521.46 -
% Change in trips - 70.05% -
Table 4: Impact of e-scooters on generation of public transport trips. Columns: (1) Mode of transport; (2) Impact of e-scooters on trips; (3) Average number of trips; (4) Percentage change in trips. Data is presented for Central, Intermediate, and Peripheral regions.
Attraction of Trips
Tables in Appendix D show the results of negative binomial models for public transport trip attraction, while Table 5 shows the marginal effects of scooter origin zones, located within different regions, on public transport trip attraction. In the Central Region, e-scooter trips are linked to a large drop in public transport trips ending in these zones. Specifically, in these central zones, e-scooter trips lead to a reduction of approximately 389.17 trips per hour, which corresponds to a 72.06% decline in bus and metro trips, given the average of 540.08 hourly trips in these areas. However, when focusing on bus trips alone, e-scooter trips attract 154.21 additional trips per hour, representing a 38.12% increase over the average of 1186.10 hourly metro trips in central zones. For metro trips, the impact of e-scooters is not statistically significant.
In zones located in the Intermediate Region, e-scooter trips positively influence the attraction of public transport trips. Each e-scooter trip generates an additional 87.17 bus or metro trips per hour, leading to a 63.85% increase over the average of 136.52 hourly trips in these intermediate zones. The impact on bus trips is more modest, with an increase of 14 additional bus trips per hour per zone, corresponding to a 22.29% rise from the region’s average of 62.78 bus trips per hour. Focusing on metro trips specifically, e-scooter trips attract approximately 286.41 trips per hour, resulting in a 62.17% increase compared to the average of 460.68 hourly metro trips in intermediate zones.
In zones located in the Peripheral Region, the impact of e-scooter trips on the attraction of public transport trips is more nuanced. Scooters only have a significant impact on bus trips, attracting 38.62 additional bus trips per hour, representing a notable 77.29% increase compared to the average of 49.97 hourly bus trips in these peripheral zones.
Mode of Transport Category Central Intermediate Peripheral
Bus or Metro E-scooter impact on trips -389.17 87.17
Average trips 540.08 136.52 -
% Change in trips -72.06% 63.85% -
Bus E-scooter impact on trips 51.16 14.00 38.62
Average trips 154.21 62.78 49.97
% Change in trips 38.12% 22.29% 77.29%
Metro E-scooter impact on trips - 286.41 -
Average trips - 460.68 -
% Change in trips - 62.17% -
Table 5: Impact of e-scooters on attraction of public transport trips. Columns: (1) Mode of transport; (2) Impact of e-scooters on trips; (3) Average number of trips; (4) Percentage change in trips. Data is presented for Central, Intermediate, and Peripheral regions.
6 Discussion
The results of this study provide a comprehensive view of the impact of e-scooters on public transportation in Santiago, Chile, considering both generations and attraction scenarios. By unifying these findings, three key aspects emerge: the regional variability of e-scooter impacts, the complementary and competitive role of e-scooters, and strategic implications for urban mobility planning.
Regional Differences
In both scenarios, we see that e-scooters affect public transportation differently in each region, based on the percentage changes per hour and per zone. In the Central Region, e-scooters compete with public transport overall, reducing combined bus and metro trips in both the generation (-21.38% per hour per zone) and attraction (-72.06% per hour per zone) scenarios. However, this region also shows a substantial increase in bus trips associated with e-scooter usage, with increases of 76.39% in trip generation and 38.12% in trip attraction, both measured per hour per zone. This pattern suggests that while e-scooters may displace some metro trips, they are facilitating access to and from buses, particularly in areas where walking distances might be longer or less convenient.
In contrast, the Intermediate Regions show predominantly complementary effects of e-scooters for both transport modes, metro and buses. In the Intermediate Region, e-scooters increase combined bus and metro usage by 56.83% in trips generation and 63.85% in trips attraction, both per hour and per zone, with a slight preference for metro usage (70.05% in generation and 62.17% in attraction). In the Peripheral Region, scooters significantly increase the generation of public transport trips, as well as the overall generation and attraction of bus trips. The impact is noticeable, with a 40.17% increase in generated trips, and in the case of bus trips, the increases are 84.64% in generation and 77.29% in attraction, all per hour and per zone. These results highlight the role of e-scooters in areas with varying levels of public transport coverage, where they can complement existing services by improving accessibility and connectivity.
By associating public transport generation trips with first-mile trips made on scooters, it is observed that scooters serve as a first-mile mode for accessing both buses and the metro in the Intermediate region, while facilitating only bus access in the Peripheral and Central regions. When considering that public transport attraction trips can also be linked to last-mile scooter trips, it becomes evident that they are used as a last-mile mode for exiting buses and the metro in the Intermediate region, and once again, only for buses in the Peripheral and Central regions.
These findings are consistent with those of Luo et al. (2021), who also found that e-scooters can either compete with or complement public transportation depending on the urban context and the availability of other transport options. However, our study goes further by considering not only buses but also the metro, providing a more comprehensive understanding of how e-scooters can be integrated into multimodal transport systems across different urban regions.
Differential Impacts on Transport Modes
An intriguing pattern emerged in our results, particularly in the Central Region. While e-scooters were associated with a decrease in combined bus and metro usage, they showed a positive impact on bus trips alone. This seemingly contradictory finding warrants further exploration. One possible explanation is that e-scooters are primarily replacing short-distance metro trips, while simultaneously improving access to bus stops for longer journeys. Additionally, the flexibility of e-scooters might allow users to avoid peak crowding times on metro services, a factor known to influence public transport choice and user experience (Tirachini et al., 2013). This could lead to an overall reduction in combined public transport trips, but an increase in bus usage specifically, as buses might be perceived as less crowded or more accessible when combined with e-scooter use.
Complementary and Competitive Role of E-Scooters
E-scooters demonstrate a dual role, acting as both complements and competitors to public transportation depending on the context. In the Central Region, with its high density and well-developed transport infrastructure, e-scooters seem to compete more directly with public transport. Specifically, e-scooters in the Central Region are associated with a reduction in combined bus and metro trips in both the generation (-21.38% per hour per zone) and attraction (-72.06% per hour per zone) scenarios. However, they also contribute to a substantial increase in bus trips, with rises of 76.39% in generation and 38.12% in attraction trips, both per hour per zone. This suggests that while e-scooters may replace public transport for certain trips, they also provide significant connections for bus users.
In contrast, in the Intermediate and Peripheral Regions, e-scooters clearly play a complementary role by connecting users to public transport and extending the reach of the urban transport system. Given the absence of metro stations in the Peripheral Regions, the results indicate that scooters effectively bridge gaps in transport accessibility, especially for bus trips. In the Intermediate Region, this pattern aligns with studies like Weschke (2023) in Berlin, which found a notable increase in e-scooter demand near metro stations, highlighting the complementary role scooters can play in connecting users to other modes of public transport.
The Peripheral Region in particular benefits from improved bus access, where the increase in public transport connectivity is crucial due to the limited transport options compared to the Central Region. This observation supports the idea that e-scooters can serve as last-mile solutions in areas with less developed transport infrastructure. In the Intermediate Region, a similar complementary pattern is observed, reinforcing e-scooters’ role in facilitating access to public transport.
While our findings highlight the complementary nature of e-scooters in both Intermediate and Peripheral Regions, they differ somewhat from studies in U.S. cities like Louisville and Nashville Ziedan et al. (2021a, b), which reported that e-scooters sometimes reduced bus demand, particularly for routine trips. However, as in our findings, these studies also acknowledged that scooters could increase bus demand for specific routes or social trips. This suggests that the relationship between e-scooters and public transport is highly context-dependent, varying by city structure, transport infrastructure, and user behavior.
Policy Implications
Investing in and promoting sustainable travel modes such as cycling and public transport is one of the most promising strategies for reducing transport-related carbon emissions (see e.g. Brand et al., 2021; Amelung et al., 2019; de Nazelle et al., 2010; Keall et al., 2018). Our study contributes to this goal by providing evidence on the effectiveness implementing e-scooters to promote the use of one of these sustainable travel modes, public transport. Specifically, the results of this study have important implications for urban mobility planning in Santiago and other cities with similar characteristics. It shows that the evaluation of e-scooter deployment should be done separately for different areas of a city. In this study we show that in peripheral and moderately dense areas, e-scooters offer an effective solution for improving access to public transport, reducing car dependency, and fostering a more integrated and efficient transport system. This study suggests to policy-makers to prioritize first the installation of e-scooters in zones with less level of accessibility to public transportation in order to incentives the use of it. In this context, policies that promote the integration of e-scooters with public transport, such as the installation of scooter stations near bus stops and metro stations, could maximize the observed benefits.
Our analysis highlights the value of using track-and-trace data (Harrison et al., 2020) from smart cards and GPS data from public transport and e-scooters to understand and influence travel behaviour. This approach enables policy makers to develop targeted, data-driven strategies to promote sustainable travel modes. Coupled with the observed variability in the impact of e-scooters across different travel contexts (different areas of the city and different travel modes), passive transport data allows the identification of specific travel contexts where the effectiveness of interventions such as e-scooter deployment is particularly high compared to other travel contexts. In this context, our study suggests that it is highly recommended to develop transport policies aimed at defining and implementing e-scooter systems targeted at commuters living in peripheral or intermediate areas.
For these strategies to be effective, it is crucial that urban planners and transport authorities have access not only to data from public transport systems but also to data from emerging modes of transport such as e-scooters. Understanding the patterns of e-scooter usage, their interaction with public transport, and the specific needs of different urban areas requires comprehensive data collection and analysis. The integration of such data into urban planning processes can help in designing more effective and responsive transport systems that cater to the diverse needs of the population, enhance multimodal connectivity, and promote sustainable urban mobility.
Limitations
Our study, while offering significant contributions to understanding the interaction between e-scooters and public transportation in Santiago, presents some limitations. First, our focus on a single urban context limits the generalizability of the results to other Latin American cities with different geographical, socioeconomic, and infrastructural characteristics. Additionally, we did not consider the impact of external factors such as seasonal changes, or extraordinary events that could influence the use of public transportation and e-scooters.
Although the use of the difference-in-differences (diff-in-diff) method improves our ability to infer causality, this approach can still be limited by unobserved factors that vary over time and could affect our results, such as changes in the quality of public transport service, fluctuations in the local economy, or variations in user safety perceptions. Additionally, our study calculates short-term effects and recognizes that the long-term complementarity of public transportation and micromobility might be underestimated as user behavior adjusts over time. Assessing their environmental impact and long-term viability as a mobility option is crucial. Future studies could benefit from additional experimental or longitudinal designs to address these limitations and provide a more robust understanding of the observed dynamics.
7 Conclusions
This study methodically explores how shared e-scooters influence public transport usage in Santiago, Chile, by analyzing trip origins and destinations. The findings indicate that e-scooters can both complement and compete with public transport, depending on the urban context and specific transport modes involved.
7.1 Summary of Key Findings
  • 1.
    Regional Variability: The impact of e-scooters on public transport differs significantly across regions. In the Central Region, e-scooters tend to compete with the metro system while complementing bus services, suggesting their role in improving access to buses in densely populated areas. In contrast, in the Intermediate and Peripheral Regions, e-scooters primarily complement public transport, enhancing accessibility and extending the reach of the transport system, particularly in areas with less developed infrastructure.
  • 2.
    Dual Roles: E-scooters serve both complementary and competitive roles in urban mobility. Their ability to either enhance or compete with public transport modes highlights the need for region-specific policies that account for these dynamics.
7.2 Strategic Implications
This investigation underscores the importance of integrating e-scooter usage data into urban mobility planning. By incorporating comprehensive data from various modes of transportation, urban planners and transport authorities can design more efficient systems that enhance multimodal connectivity and address diverse population needs.
7.3 Implications for Policy and Practice
  • 1.
    Data Integration: Urban mobility strategies should include data from both traditional public transport systems and new transportation modes like e-scooters. This integration is essential for understanding usage patterns, optimizing transport networks, and fostering sustainable urban mobility.
  • 2.
    Tailored Strategies: In Santiago de Chile, region-specific strategies are needed to manage the role of e-scooters in urban transport. In central areas, e-scooter policies can focus on enhancing their role alongside bus services. In intermediate regions, fostering synergy between e-scooters, metro, and bus networks can improve transport efficiency. In peripheral regions, promoting e-scooter use, particularly in combination with bus services, can enhance transport accessibility. For other cities, it is important to study the specific configuration and distribution of their transport systems to adapt these strategies to local contexts.
By addressing these strategic areas, cities like Santiago de Chile, as well as others, can develop more nuanced and effective urban mobility frameworks that balance the roles of e-scooters and public transport.
Acknowledgments
We thank Gonzalo Cortés from Scoot Networks for facilitating the data for this study. D. Opitz and M. Rivas were partially funded by ANID Fondecyt de Iniciación #11220799. E. Graells-Garrido was partially funded by National Center for Artificial Intelligence CENIA FB210017, Basal ANID. N. Meza was partially funded by FCFM starter project “Modelo Eficiente de Fusión de Datos Urbanos.”
Appendix A Time blocks
Time Period
06:01 - 09:00:59 Morning Peak
09:01 - 12:00:59 Morning Valley
12:01 - 14:00:59 Lunch
14:01 - 17:30:59 Afternoon Valley
17:31 - 20:30:59 Afternoon Peak
20:31 - 23:00:59 Night Valley
23:01 - 06:00:59 Night
Table 6: Time blocks used in the study
Appendix B Supplementary Summary of Zone Clustering Variables
Name min max
under age population 3.25 39.35
working population 56.28 93.30
retired population 1.22 24.84
female population 10.90 56.44
avg. educational level 8.32 15.98
bus stops 0.00 80.00
bus stop density 0.00 48.25
bus stops per 1000 people 0.00 77.93
population density 31.88 34993.67
avg. number of trips per person 1.00 2.71
avg. time of first trip 9.49 13.05
avg. trip distance 1627.34 20322.69
Table 7: Summary of Zone Clustering Variables. This table provides the minimum and maximum values for each variable used in our study.
Appendix C Results of Negative Binomial Models for Evaluating Generation of Public Transport Trips
Parameters Peripheral Intermediate Central
Region Region Region
Intercept 3.6948 (92.957) *** 2.9166 (58.573) *** 3.1445 (21.757) ***
Time periods
Morning peak 0.5638 (41.678) *** 0.2558 (18.403) *** -0.0793 (-2.284) *
Lunch -0.1860 (-12.482) *** -0.0128 (-0.843) 0.1897 (5.074) ***
Afternoon peak -0.1729 (-13.114) *** 0.1939 (14.360) *** 0.5787 (16.969) ***
Afternoon valley -0.1559 (-11.882) *** 0.1558 (11.573) *** 0.4729 (13.886) ***
Night -1.7074 (-118.555) *** -2.3861 (-169.260) *** -2.7928 (-86.563) ***
Night valley -1.1541 (-81.095) *** -0.5905 (-41.059) *** -0.3549 (-10.062) ***
Type of day
Friday 0.0435 (4.034) *** 0.0449 (4.096) *** 0.0446 (1.669)
Saturday -0.4696 (-42.194) *** -0.4781 (-42.562) *** -0.5387 (-19.821) ***
Sunday -0.8738 (-75.446) *** -0.9108 (-78.426) *** -1.0786 (-38.667) ***
Treatment period
Year 2019 0.0166 (2.231) * -0.0470 (-6.107) *** 0.0200 (1.087)
Treatment zones
Scooter trips destination 0.4079 (3.634) *** 0.5880 (25.502) *** -0.2258 (-2.867) **
Infrastructure
Metro 0.1310 (16.203) *** 0.0170 (2.169) * 0.1969 (5.786) ***
Sociodemographic information
Log(working population) 0.1564 (35.042) *** 0.2813 (47.675) *** 0.3028 (18.138) ***
Relative educational level 0.1692 (34.307) *** 0.2125 (44.320) *** 0.5689 (44.883) ***
Model performance
Nº observations 103,475 119,855 21,944
Log-likelihood -5.6167e+05 -6.8337e+05 -1.4619e+05
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 8: Results of negative binomial models for evaluating generation of bus and metro trips
Parameters Peripheral Intermediate Central
Region Region Region
Morning peak 62.753 (1.580) *** 43.412 (2.378) *** -47.985 (21.022) **
Lunch -20.707 (1.665) *** -2.164 (2.568) 114.830 (22.726) ***
Afternoon peak -19.248 (1.474) *** 32.898 (2.300) *** 350.333 (21.615) ***
Afternoon valley -17.353 (1.465) *** 26.432 (2.291) *** 286.268 (21.268) ***
Night -190.052 (1.954) *** -404.928 (3.278) *** -1690.551 (30.204) ***
Night valley -128.461 (1.766) *** -100.212 (2.528) *** -214.808 (21.662) ***
Year 2019 1.846 (0.827) ** -7.982 (1.308) *** 12.130 (11.154)
Friday 4.843 (1.201) *** 7.624 (1.862) *** 27.000 (16.188)
Saturday -52.273 (1.281) *** -81.128 (1.973) *** -326.066 (17.207) ***
Sunday -97.262 (1.414) *** -154.568 (2.182) *** -652.903 (19.453) ***
Scooter trips destinations 45.404 (12.498) *** 99.799 (4.007) *** -136.663 (47.625) ***
Metro 14.584 (0.904) *** 2.883 (1.329) ** 119.174 (20.697) ***
Log(working population) 17.410 (0.507) *** 47.739 (1.309) *** 183.321 (10.685) ***
Relative educational level 18.830 (0.566) *** 36.069 (0.852) *** 344.369 (9.459) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 9: Marginal effects of negative binomial models for evaluating generation of bus and metro trips
Parameters Peripheral Intermediate Central
Region Region Region
Intercept 2.387 (67.912) *** 1.565 (36.026) *** 3.207 (27.911) ***
Time periods
Morning peak 0.553 (45.315) *** 0.229 (19.929) *** 0.003 (0.087)
Lunch -0.160 (-11.900) *** 0.018 (1.401) 0.166 (5.277) ***
Afternoon peak -0.142 (-11.869) *** 0.173 (15.519) *** 0.487 (17.079) ***
Afternoon valley -0.118 (-10.011) *** 0.173 (15.585) *** 0.438 (15.381) ***
Night -1.510 (-116.152) *** -1.915 (-162.572) *** -1.955 (-71.144) ***
Night valley -1.144 (-88.904) *** -0.584 (-48.888) *** -0.434 (-14.533) ***
Type of day
Friday 0.034 (3.518) *** 0.030 (3.337) *** 0.036 (1.621)
Saturday -0.478 (-47.719) *** -0.481 (-51.589) *** -0.475 (-20.657) ***
Sunday -0.879 (-83.864) *** -0.898 (-93.109) *** -0.937 (-39.569) ***
Treatment period
Year 2019 -0.022 (-3.376) *** -0.056 (-8.641) *** -0.206 (-13.259) ***
Treatment zones
Scooter trips destination 0.866 (8.604) *** 0.229 (12.127) *** 0.781 (11.980) ***
Infrastructure
Metro 0.143 (19.893) *** -0.091 (-13.866) *** 0.491 (16.917) ***
Sociodemographic information
Log(working population) 0.273 (70.063) *** 0.379 (72.816) *** 0.170 (13.165) ***
Relative educational level 0.105 (25.776) *** 0.145 (37.522) *** 0.313 (28.425) ***
Model performance
Nº observations 102,881 118,438 21,486
Log-Likelihodd -5.3519e+05 -6.1882e+05 -1.2555e+05
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 10: Results of negative binomial models for evaluating generation of bus trips
Parameters Peripheral Intermediate Central
Region Region Region
Morning peak 46.4909 (1.073)*** 20.1829 (1.019)*** 0.4706 (5.425)
Lunch -13.4742 (1.135)*** 1.5496 (1.106) 30.8586 (5.866)***
Afternoon peak -11.9116 (1.006)*** 15.2159 (0.984)*** 90.5389 (5.473)***
Afternoon valley -9.9645 (0.997)*** 15.2425 (0.982)*** 81.4608 (5.440)***
Night -127.0872 (1.289)*** -168.4752 (1.298)*** -363.5232 (6.644)***
Night valley -96.2803 (1.201)*** -51.3411 (1.084)*** -80.6726 (5.657)***
Year 2019 -1.8997 (0.563)*** -4.8398 (0.561)*** -38.3488 (2.944)***
Friday 2.8771 (0.818)*** 2.6661 (0.799)** 6.7609 (4.173)
Saturday -40.2679 (0.874)*** -42.2632 (0.848)*** -88.2394 (4.411)***
Sunday -73.7713 (0.966)*** -78.9807 (0.936)*** -174.2467 (4.877)***
Scooter trips destinations 72.8203 (8.481)*** 20.1006 (1.663)*** 145.2156 (12.422)***
Metro 12.0656 (0.610)*** -8.0391 (0.582)*** 91.3796 (5.520)***
log(Working population) 23.0446 (0.350)*** 33.3239 (0.492)*** 31.7262 (2.466)***
Relative educational level 8.8441 (0.347)*** 12.7893 (0.349)*** 58.2051 (2.209)***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 11: Marginal effects of negative binomial models for evaluating generation of bus trips
Parameters Intermediate Central
Region Region
Intercept 7.724 (66.725) *** 4.896 (21.553) ***
Time periods
Morning peak 0.350 (11.775) *** -0.114 (-2.537) **
Lunch -0.100 (-3.164) *** 0.202 (4.242) ***
Afternoon peak 0.323 (11.080) *** 0.672 (15.188) ***
Afternoon valley 0.216 (7.450) *** 0.535 (12.125) ***
Night -2.831 (-79.758) *** -2.950 (-56.509) ***
Night valley -0.566 (-18.605) *** -0.306 (-6.834) ***
Type of day
Friday 0.066 (2.758) *** 0.042 (1.159)
Saturday -0.478 (-18.935) *** -0.653 (-17.247) ***
Sunday -0.899 (-33.717) *** -1.305 (-33.075) ***
Treatment period
Year 2019 -0.091 (-5.187) *** -0.043 (-1.696)
Treatment zones
Scooter trips destination 0.714 (17.166) *** 0.385 (1.456)
Infrastructure
Metro
Sociodemographic information
Log(working population) -0.182 (-13.377) *** 0.202 (7.693) ***
Relative educational level 0.255 (23.487) *** 0.411 (21.753) ***
Model performance
Nº observations 20,079 7,996
Log-likelihood -1.3852e+05 -61517
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 12: Results of negative binomial model for evaluating generation of metro trips
Parameters Intermediate Central
Region Region
Intercept -93.041 (-13.088) *** 183.3205 (17.157) ***
Morning peak 178.7567 (11.575) *** -47.9854 (-2.283) **
Lunch -50.9498 (-3.161) ** 114.8304 (5.053) ***
Afternoon peak 165.2241 (10.947) *** 350.3334 (16.208) ***
Afternoon valley 110.6648 (7.406) *** 286.2683 (13.460) ***
Night -1447.3674 (-57.926) *** -1690.5511 (-55.971) ***
Night valley -289.2255 (-17.965) *** -214.8078 (-9.916) ***
Friday 33.9532 (2.748) ** 27.0006 (1.668) *
Saturday -244.4253 (-18.360) *** -326.0656 (-18.950) ***
Sunday -459.9325 (-30.774) *** -652.9032 (-33.563) ***
Year 2019 -46.5152 (-5.174) *** 12.1303 (1.087)
Scooter trip destination 365.3049 (16.333) *** -136.6634 (-2.870) **
Metro -93.0410 (-13.088) *** 119.1737 (5.758) ***
Log(working population) 183.3205 (17.157) *** 183.3205 (17.157) ***
Relative educational level 130.5021 (22.234) *** 344.3689 (36.405) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 13: Marginal effects of negative binomial models for evaluating generation of metro trips
Appendix D Results of Negative Binomial Models for Evaluating Attraction of Public Transport Trips
Parameters Peripheral Intermediate Central
Region Region Region
Intercept 3.363 (88.808) *** 2.643 (57.671) *** 3.172 (21.786) ***
Time periods
Morning peak 0.284 (21.874) *** 0.148 (11.308) *** 0.159 (4.707) ***
Lunch 0.221 (16.331) *** 0.015 (1.097) -0.166 (-4.604) ***
Afternoon peak 0.996 (79.981) *** 0.409 (32.227) *** -0.081 (-2.454) *
Afternoon valley 0.631 (51.610) *** 0.267 (21.396) *** -0.043 (-1.313)
Night -1.420 (-100.531) *** -2.207 (-165.063) *** -3.025 (-96.222) ***
Night valley 0.328 (24.474) *** -0.378 (-27.806) *** -1.019 (-28.876) ***
Type of day
Friday 0.024 (2.346) ** 0.038 (3.735) *** 0.041 (1.611)
Saturday -0.482 (-45.596) *** -0.488 (-46.464) *** -0.533 (-20.183) ***
Sunday -0.919 (-83.369) *** -0.930 (-85.292) *** -1.071 (-39.088) ***
Treatment period
Year 2019 0.010 (1.408) -0.046 (-6.388) *** -0.085 (-2.714) ***
Treatment zones
Scooter trips origins 0.188 (1.425) 0.654 (26.747) *** -0.762 (-9.003) ***
Infrastructure
Metro 0.179 (23.061) *** 0.065 (8.970) *** 0.121 (3.693) ***
Sociodemographic information
Log(working population) 0.075 (17.477) *** 0.268 (48.505) *** 0.317 (13.165) ***
Relative educational level 0.165 (34.366) *** 0.278 (63.178) *** 0.602 (49.025) ***
Model performance
Nº observations 130,832 152,443 27,018
Log-likelihood -6.5106e+05 -8.2575e+05 -1.7275e+05
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 14: Results of negative binomial models for evaluating attraction of metro and bus trips
Parameters Peripheral Intermediate Central
Region Region Region
Morning peak 19.716 (1.745) *** 81.130 (17.283) *** 20.006 (0.920) ***
Lunch 2.019 (1.840) -84.628 (18.433) *** 15.592 (0.959) ***
Afternoon peak 54.549 (1.736) *** -41.443 (16.904) ** 70.270 (0.991) ***
Afternoon valley 35.610 (1.684) *** -21.880 (16.662) 44.500 (0.906) ***
Night -294.399 (2.366) *** -1544.635 (26.621) *** -100.253 (1.115) ***
Night valley -50.349 (1.833) *** -520.130 (19.738) *** 23.117 (0.955) ***
Year 2019 -6.067 (0.950) *** 24.765 (9.131) ** 0.699 (0.496)
Friday 5.067 (1.356) *** 21.203 (13.166) 1.678 (0.715) **
Saturday -65.059 (1.453) *** -272.559 (14.120) *** -34.027 (0.772) ***
Sunday -124.000 (1.623) *** -547.100 (16.052) *** -64.856 (0.861) ***
Scooter trips origins 87.171 (3.346) *** -389.172 (43.670) *** 13.275 (9.314)
Metro 8.695 (0.971) *** 61.877 (16.794) *** 5.319 (0.305) ***
Log(working population) 35.683 (0.759) *** 161.874 (9.187) *** 11.659 (0.349) ***
Relative educational level 37.051 (0.638) *** 307.614 (7.839) *** 11.659 (0.349) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 15: Marginal effects of negative binomial models for evaluating attraction of bus and metro trips
Parameters Peripheral Intermediate Central
Region Region Region
Intercept 1.595 (48.215) *** 1.339 (34.502) *** 3.336 (31.032) ***
Time periods
Morning peak 0.291 (25.449) *** 0.141 (13.393) *** 0.155 (5.860) ***
Lunch 0.192 (16.077) *** -0.009 (-0.777) -0.170 (-5.971) ***
Afternoon peak 0.866 (78.818) *** 0.292 (28.503) *** -0.211 (-8.137) ***
Afternoon valley 0.561 (51.997) *** 0.215 (21.315) *** -0.046 (-1.803)
Night -1.271 (-101.055) *** -1.818 (-166.492) *** -2.199 (-88.225) ***
Night valley 0.234 (19.854) *** -0.438 (-39.816) *** -1.180 (-42.493) ***
Type of day
Friday 0.008 (0.918) 0.027 (3.261) *** 0.034 (1.620)
Saturday -0.477 (-51.174) *** -0.453 (-53.279) *** -0.434 (-20.779) ***
Sunday -0.875 (-89.852) *** -0.859 (-97.042) *** -0.887 (-40.966) ***
Treatment period
Year 2019 -0.036 (-5.853) *** -0.065 (-11.305) *** -0.187 (-13.207) ***
Treatment zones
Scooter trips origins 0.789 (5.829) *** 0.225 (11.543) *** 0.391 (5.897) ***
Infrastructure
Metro 0.144 (21.445) *** -0.040 (-6.774) *** 0.418 (15.787) ***
Sociodemographic information
Log(working population) 0.246 (55.428) *** 0.354 (75.494) *** 0.157 (12.710) ***
Relative educational level 0.060 (15.574) *** 0.184 (63.178) *** 0.336 (33.561) ***
Model performance
Nº observations 130,168 151,127 26,685
Log-likelihood -6.1337e+05 -7.4013e+05 -1.4669e+05
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 16: Results of negative binomial models for evaluating attraction of bus trips
Parameters Peripheral Intermediate Central
Region Region Region
Morning peak 14.253 (0.564) *** 8.779 (0.657) *** 20.363 (3.485) ***
Lunch 9.403 (0.587) *** -0.541 (0.695) -22.236 (3.736) ***
Afternoon peak 42.421 (0.589) *** 18.159 (0.645) *** -27.713 (3.426) ***
Afternoon valley 27.473 (0.549) *** 13.374 (0.632) *** -6.057 (3.360)
Night -62.248 (0.674) *** -113.066 (0.832) *** -288.239 (4.559) ***
Night valley 11.502 (0.582) *** -27.219 (0.696) *** -154.566 (4.074) ***
Year 2019 -1.772 (0.303) *** -4.051 (0.359) *** -24.431 (1.878) ***
Friday 0.401 (0.437) 1.678 (0.512) *** 4.383 (2.660)
Saturday -23.383 (0.472) *** -28.196 (0.546) *** -56.828 (2.811) ***
Sunday -42.875 (0.523) *** -53.444 (0.603) *** -116.202 (3.112) ***
Scooter trips origins 38.622 (5.662) *** 13.996 (1.217) *** 51.162 (8.711) ***
Metro 7.073 (0.332) *** -2.502 (0.370) *** 54.786 (3.524) ***
Log(working population) 12.062 (1.492) *** 22.048 (0.311) *** 12.062 (1.644) ***
Relative educational level 2.930 (0.189) *** 11.467 (0.224) *** 44.086 (1.422) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 17: Marginal effects of negative binomial models for evaluating attraction of bus trips
Parameters Intermediate Central
Region Region
Time periods
Morning peak 0.262 (9.040) *** 0.224 (4.848) ***
Lunch 0.033 (1.072) -0.174 (-3.577) ***
Afternoon peak 0.543 (19.337) *** -0.038 (-0.849)
Afternoon valley 0.350 (12.652) *** -0.025 (-0.565)
Night -2.371 (-71.668) *** -3.070 (-60.407) ***
Night valley -0.319 (-10.749) *** -0.985 (-20.666) ***
Treatment period
Year 2019 -0.061 (-4.076) *** 0.045 (1.786)
Type of day
Friday 0.066 (2.468) ** 0.044 (1.198)
Saturday -0.553 (-23.068) *** -0.637 (-16.603) ***
Sunday -0.978 (-38.578) *** -1.173 (-28.528) ***
Treatment zones
Scooter trips origins 0.627 (14.874) *** 0.395 (1.408)
Sociodemographic information
Log(working population) -0.175 (-13.360) *** 0.089 (3.253) ***
Relative educational level 0.329 (32.655) *** 0.482 (25.921) ***
Model performance
Nº observations 24,579 9,283
Log-likelihood -1.6604e+05 -70503
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 18: Results of negative binomial models for evaluating attraction of metro trips
Parameters Intermediate Central
Region Region
Morning peak 119.811 (13.318) *** 259.532 (53.871) ***
Lunch 14.836 (13.838) -201.826 (56.595) ***
Afternoon peak 247.817 (13.417) *** -44.194 (52.047)
Afternoon valley 159.727 (12.879) *** -29.187 (51.691)
Night -1082.687 (19.720) *** -3560.477 (89.221) ***
Night valley -145.858 (13.658) *** -1142.470 (60.046) ***
Year 2019 -30.495 (7.489) *** 52.561 (29.450)
Friday 25.798 (10.462) ** 50.876 (42.487)
Saturday -252.478 (11.465) *** -738.275 (47.236) ***
Sunday -446.582 (12.996) *** -1361.016 (55.188) ***
Scooter trips origins 286.414 (19.966) *** 458.374 (325.778)
Log(working population) -79.894 (6.110) *** 102.957 (31.733) ***
Relative educational level 150.112 (5.106) *** 559.393 (24.861) ***
*p $<$ 0.10; **p $<$ 0.05; ***p $<$ 0.01. Z-values are shown in parenthesis.
Table 19: Marginal effects of negative binomial models for evaluating attraction of metro trips